### v1

# The Algorithmic Regulator

Abstract
The regulator theorem states that, under certain conditions, any optimal controller must embody a model of the system it regulates, grounding the idea that controllers embed, explicitly or implicitly, internal models of the controlled. This principle underpins neuroscience and predictive brain theories like the Free-Energy Principle or Kolmogorov/Algorithmic Agent theory. However, the theorem is only proven in limited settings. Here, we treat the deterministic, closed, coupled world-regulator system ($W$,$R$) as a single self-delimiting program $p$ via a constant-size wrapper that produces the world output string $x$ fed to the regulator. We analyze regulation from the viewpoint of the algorithmic complexity of the output, $K(x)$. Intuitively, knowing that $K(x)$ is small under regulation prunes the class of minimal candidate worlds, but this cannot, in general, certify mutual algorithmic information of world and regulator. This motivates a probabilistic approach. We define $R$ to be a good algorithmic regulator if it reduces the algorithmic complexity of the readout relative to a null (unregulated) baseline $\varnothing$, i.e., $\Delta=K\big(O_{W,\varnothing}\big)-K\big(O_{W,R}\big)>0.$ We then prove that the larger $\Delta$ is, the more world-regulator pairs with high mutual algorithmic information are favored. More precisely, a complexity gap $\Delta>0$ yields $\Pr((W,R)\mid x)\leq C\,2^{\,M(W{:}R)}\,2^{-\Delta}$, making low M(W:R)M(W{:}R) exponentially unlikely as $\Delta$ grows. This is an AIT version of the idea that “the regulator contains a model of the world.” The framework is distribution‑free, applies to individual sequences, and complements the Internal Model Principle. Beyond this necessity claim, the same coding‑theorem calculus singles out a canonical scalar objective and implicates a planner. On the realized episode, a regulator behaves as if it minimized the conditional description length of the readout.
1 Introduction
In the Kolmogorov Theory (KT) of consciousness, an algorithmic agent is a system that maintains (tele)homeostasis (persistence of self or kind) by learning and running succinct generative models of its world coupled to an objective function and a action planner [26, 28, 27]. Closely related, Active Inference (AIF) models biological agents as minimizing variational free energy under a generative model [15, 24]. These frameworks suggest that “agents with world‑modeling engines, objective functions, and planners” are natural minimal models of homeostasis (goal‑conditioned setpoint control). But for the kinds of homeostatic systems we actually encounter in nature (cells, organisms, engineered servos), how can we tell—operationally—whether they are algorithmic agents in this sense?
The classical cybernetics statement that “every good regulator of a system must be a model of that system” originates with Conant and Ashby’s 1970 paper (the Good Regulator Theorem, GRT) [8]. While influential, the GRT has been criticized for the looseness of its definitions of “model” and “goodness”, and for a proof that does not clearly deliver the headline claim  [5]. In modern control theory, the rigorous statement that fills a similar conceptual niche is the Internal Model Principle (IMP): under appropriate hypotheses, perfect regulation or disturbance rejection for a given signal class requires that the controller embed a dynamical copy of the signal generator [12, 13, 34]. The IMP is precise (and falsifiable) within its scope, and is now a standard backbone for robust control; see [6] for a contemporary review across control, bioengineering, and neuroscience. However, the classical IMP is a linear result: for finite-dimensional LTI plants ( linear, time-invariant meaning the system matrices do not change with time) and exogenous signals generated by a finite-dimensional LTI exosystem, robust asymptotic tracking/disturbance rejection requires that the controller embed a copy of the exosystem dynamics [14]. For nonlinear systems, the appropriate generalization is the nonlinear output-regulation framework: if the regulator equations admit smooth solutions and the plant’s zero dynamics on the regulated manifold are (locally) stable, together with suitable immersion/detectability assumptions, then one can construct dynamic output-feedback regulators that embed a (possibly adaptive) internal model and achieve local or semiglobal robust regulation [20, 17, 25]. However, absent these structural hypotheses, a complete nonlinear analogue of the IMP with the same necessity/robustness guarantees as in the LTI case is not generally available. Table 2 provides a comparison of the different regulator theorem statements, which can be compared with the one presented here.
In this paper, we recast the modeling requirement in a setting independent of linearity, probability, or specific signal classes, by using algorithmic information theory (AIT). We model a world $W$ and a regulator $R$ as deterministic causal Turing machines that interact over interface tapes. We denote the world output by $x=O_{W}$ (over some temporal horizon of length $N$). Our main technical claim is that regulation in the algorithmic sense, i.e., simplicity, forces algorithmic dependence between $W$ and $R$.
A model in the present context is a program capable of compressing (or generating) data. Similarly, “the regulator contains a model of the world” is interpreted in an algorithmic‑information sense: the regulator $R$ carries nontrivial information about $W$, quantified by positive mutual algorithmic information M(W:R)>0M(W{:}R)>0 (up to the standard $O(\log)$ slack). Equivalently, knowing $R$ makes the shortest description of $W$ strictly shorter, $K(W\mid R)<K(W)$. This notion does not require $R$ to embed a dynamical copy of $W$; rather, it formalizes “model content” as mutual algorithmic information.
We formalize this with the following definition:
Definition 1.1 (Algorithmic “internal model”).
Given a fixed horizon $N$ (implicitly conditioned), we say that $R$ contains an internal model of $W$ in the algorithmic sense if M(W:R)>0M(W{:}R)>0 (up to $O(\log)$), equivalently $K(W\mid R)<K(W)$. The magnitude of M(W:R)M(W{:}R) quantifies the amount of computable structure in $W$ that $R$ carries.
The definition ground on mutual algorithmic information $M$ is further motivated by the following: (i) Machine invariance: $M$ is invariant up to $O(1)$ under changes of universal machine. (ii) Distribution‑free: $M$ is defined for individual objects (programs), not probabilistic models. (iii) Operational meaning: M(W:R)M(W{:}R) is precisely the codelength reduction in describing $W$ when $R$ is known, aligning with MDL/Occam reasoning via the Coding Theorem [33, 44, 22]. (iv) Complement to IMP: Unlike the IMP—which requires a dynamical replica of the exosystem under structural assumptions—M(W:R)M(W{:}R) captures any computable shared structure evidenced by contrastive compressibility.
This is the appropriate lens for our contrastive results, and it complements the Internal Model Principle, where “model” means a dynamical replica of the Exosystem (a part of the World in our framework, see Figure 2) under stated structural hypotheses [14, 20, 17, 25]. Conceptually, our AIT result is complementary to the IMP: whereas the IMP states what structural content must be present in a controller to achieve perfect regulation for a given signal class [12, 13, 34], our results quantify how much algorithmic information the regulator must carry about the world whenever it succeeds in making the measured outcome compressible.
We score regulation by how compressible a task-weighted readout is. Let $x_{t}$ be the weighted error, and let $x_{1:T}$ denote the $T$-sample string. Fix a prefix-free lossless code (e.g., a universal compressor), and define the per-sample code length $L_{T}:=\tfrac{1}{T}\,L_{C}(x_{1:T})$. A regulator $R$ is better on horizon $T$ when it makes $L_{T}$ smaller than a null baseline $\varnothing$, i.e., when the contrastive gap $\Delta:=L_{T}(x;\,\varnothing)-L_{T}(x;\,R)$ is positive. This choice is natural: for stationary ergodic data, the normalized code length of any universal compressor converges to the Shannon entropy rate $h(x)$: one has (almost surely) $\lim_{T\to\infty}K(x_{1:T})/T=h(x)$ [9, 22], so the Kolmogorov‑based criterion reduces to the Shannon criterion whenever those stochastic assumptions are appropriate—while remaining meaningful outside them. For Gaussian processes,
$h(x)\;=\;\frac{1}{4\pi}\int_{-\pi}^{\pi}\log\!\big(2\pi e\,S_{xx}(\omega)\big)\,d\omega,$
so attenuating the in‑band sensitivity (reducing $S_{xx}$ where it matters [42, 16]) reduces codelength. In short, “compressible error” coincides with the classical intuition that good regulation removes variability/uncertainty in the task band and the IMP [4, 14].
In the next sections, we first provide an overview of the AIT setting and of the results, followed by the analysis of the single episode scenario. The next section provides a formal definition of the algorithmic regulator and the corresponding theorem.
2 Setting
Unless stated otherwise, $U$ is the standard three–tape universal prefix Turing machine: a read‑only input tape holding a self‑delimiting program $p$, a work tape (private scratch memory), and a write‑only output tape. When we write $U(p)=x$ we mean that, upon halting, the contents of the output tape equal $x$; the work tape is never part of the scored output. The domain of halting programs is prefix‑free, so Kraft–McMillan applies and the universal a priori semimeasure $m(x)=\sum_{U(p)=x}2^{-|p|}$ is well defined. By the invariance theorem, replacing $U$ by any other universal prefix machine (single‑ or multi‑tape) changes all complexities only by an additive $O(1)$; all Coding‑Theorem statements we use depend only on prefix‑freeness and therefore remain valid up to these constants (see, e.g., [22]).
The (prefix) Kolmogorov complexity of $x$ is the length of its shortest description,
$K(x)\ :=\ \min\{\,|p|:\ U(p)=x\,\}.$
Intuitively, $K(x)$ is the best achievable compressed size of $x$ on $U$. If $K(x)\ll|x|$, then $x$ has a short generative regularity; if $K(x)\approx|x|$, $x$ is (algorithmically) random. By the invariance theorem, $K$ is machine‑independent up to an additive constant $O(1)$ [22]. A fundamental limitation is that $K(\cdot)$ is not computable: no algorithm can output $K(x)$ for all $x$ [7, 22]. However, algorithms for upper bounds of $K(x)$ exist, as we discuss below.
Given auxiliary data $y$ on a read‑only auxiliary tape, the conditional complexity
$K(x\mid y)\ :=\ \min\{\,|p|:\ U(p,y)=x\,\}$
is the shortest description of $x$ given $y$. It operationalizes how much new information is needed to reconstruct $x$ once $y$ is known (e.g., “world given regulator,” or “output given model”).
The mutual algorithmic information (up to the usual $O(\log)$ slack) is
[…]
M(x:y)M(x{:}y) measures the algorithmically shared structure between $x$ and $y$: how many bits we save when describing one with the help of the other. In our setting, “the regulator contains a model of the world” means M(W:R)>0M(W{:}R)>0 (information‑theoretic dependence), not necessarily a dynamical replica.
Intuitively, strings produced by shorter programs are more likely. Solomonoff–Levin’s universal a priori semimeasure $m(x)$ and the Coding Theorem link probability and description length:
$\log_{2}m(x)\ =\ K(x)\ \pm O(1),$ (1)
providing a universal Occam calculus over individual strings. [33, 32, 43, 22].
In what follows, a finite temporal horizon $N$ is fixed throughout; unless stated otherwise, we implicitly condition on $N$ (e.g., write $K(x)$ for $K(x\mid N)$). All $O(1)$ constants depend only on the choice of $U$ (and the fixed constant‑overhead wrapper that decodes $(W,R)$ and simulates their coupling to print the readout), never on particular strings; see Appendix A.2.
2.1 The Coupled World-Regulator System
We work with 3-tape Turing machines $W$ and $R$ (see Figure 1 and Appendix A.2). We identify each machine with its minimal self–delimiting program ($|W|=K(W)$, $|R|=K(R)$) [22]. A horizon $N\in\mathbb{N}$ is fixed and all complexities are conditioned on $N$ unless otherwise stated. $W$ and $R$ interact causally for $N$ steps, producing a deterministic readout $O^{(N)}_{W,R}\in\{0,1\}^{N}$. The dynamical equations are
$O_{W}=W(O_{R}),\>O_{R}=R(O_{W}).$ (2)
The performance of the regulator is evaluated from the complexity of the output, $K(x)$. Intuitively, a good regulator produces outputs of lower complexity than the unregulated case. Since $x=O^{(N)}_{W,R}$ is computable from $(W,R,N)$,
$K(x)\;\leq\;K(W,R)+O(1)\;=\;K(W)+K(R)-M(W\!:\!R)+O(1).$ (3)
Figure 1: Regulation scenario. A) A good regulator $R$ interacts with the world $W$ so that the readout $x=O_{W}$ of the world’s output is clamped to a simple, highly compressible sequence (e.g., almost all zeros). B) When the regulator is turned off, the output is more complex.
To disentangle the role of $R$ from the coarse event $K(O^{(N)})$ is small,” we fix a null regulator $\varnothing$ (where $R$’s output is set to zero). We compare the events
$E^{R}_{a}:\ K(O^{(N)}_{W,R})=a\quad\text{vs}\quad E^{\varnothing}_{b}:\ K(O^{(N)}_{W,\varnothing})=b,$ (4)
with $b>a$. Event $E^{\varnothing}_{b}$ rules out worlds that produce a simple output without regulation; the intersection $E^{R}_{a}\wedge E^{\varnothing}_{b}$ isolates $R$’s contribution.
For notational simplicity, the on–case and off–case readouts are also expressed as
$x:=O^{(N)}_{W,R},\qquad y:=O^{(N)}_{W,\varnothing}$.
For a fixed time horizon, we write $O_{W}$ for the full output produced by $W$ when coupled to $R$.
In the next sections, we provide our main results regarding mutual information between world and regulator, and implications for inferring agent-like behavior in the regulator.
3 Probabilistic Regulator Theorems
3.1 Posterior form, given the observed $x$
Lemma 3.1 (Program posterior given $x$).
With prefix prior $P(p)=2^{-|p|}$ and deterministic likelihood $P(x\mid p)=\mathbf{1}\{U(p)=x\}$,
$P(p\mid x)=\frac{2^{-|p|}}{m(x)}.$
Consequently, by (5),
$\frac{1}{c_{2}}\,2^{\,K(x)-|p|}\ \leq\ P(p\mid x)\ \leq\ \frac{1}{c_{1}}\,2^{\,K(x)-|p|}.$
Proof.
For any finite string $x$,
$K(x):=\min\{|p|:U(p)=x\},\qquad m(x):=\sum_{p:\,U(p)=x}2^{-|p|}.$
(recall Eq. 1). The Coding Theorem gives machine-dependent constants $c_{1},c_{2}>0$ with
$c_{1}\,2^{-K(x)}\ \leq\ m(x)\ \leq\ c_{2}\,2^{-K(x)}.$ (5)
Now, briefly, Bayes’ rule yields $\Pr\{p\mid x\}=2^{-|p|}/m(x)$; apply (5). In more detail, place the prefix prior $P(p)=2^{-|p|}$ on programs $p$ and use the deterministic likelihood $P(x\mid p)=\mathbf{1}\{U(p)=x\}$. Then the evidence is $P(x)=m(x)$ and the posterior is
$P(p\mid x)\;=\;\frac{P(x\mid p)P(p)}{P(x)}\;=\;\begin{cases}\displaystyle\frac{2^{-|p|}}{m(x)},&U(p)=x,\\[6.0pt] 0,&\text{otherwise.}\end{cases}$
Then, for any $p$ with $U(p)=x$,
$\frac{1}{c_{2}}\,2^{K(x)-|p|}\;\leq\;P(p\mid x)\;=\;\frac{2^{-|p|}}{m(x)}\;\leq\;\frac{1}{c_{1}}\,2^{K(x)-|p|}.$
The relation between $K(x)$ and $m(x)$ holds only up to an additive $O(1)$ term in $K$, which becomes a multiplicative constant on $m(x)$. This $O(1)$ ambiguity is unavoidable and depends on the choice of universal prefix machine $U$; $c_{1},c_{2}$ absorb exactly this machine-dependent slack. ∎
Now, in our setting the world $W$ and regulator $R$ are programs that interact for $N$ steps, producing the on-case readout $x:=O^{(N)}_{W,R}$. A fixed, constant-overhead wrapper decodes a shortest description of $(W,R)$ and simulates the coupling to print $x$ (decode + simulate); if $p_{W,R}$ denotes this canonical code, then
$|p_{W,R}|=K(W,R)+O(1),\qquad P\big((W,R)\mid x\big)\ \in\ \Big[\tfrac{1}{\tilde{c}_{2}},\tfrac{1}{\tilde{c}_{1}}\Big]\cdot 2^{\,K(x)-K(W,R)},$ (6)
for constants $\tilde{c}_{i}:=2^{O(1)}c_{i}$.
Now we can use the definition of mutual algorithmic information (up to the usual $O(\log)$ slack) to write
M(W:R)=K(W)+K(R)K(W,R)M(W{:}R)\ =\ K(W)+K(R)-K(W,R)\
and derive our first result:
Theorem 3.1.
$P\big((W,R)\mid x\big)\ \in\ \Big[\tfrac{1}{\tilde{c}_{2}},\tfrac{1}{\tilde{c}_{1}}\Big]\cdot 2^{\,K(x)-K(W)-K(R)+M(W{:}R)}<\frac{1}{\tilde{c}}\,2^{M(W{:}R)}$ (7)
3.2 The Good Algorithmic Regulator and Posterior with Contrast
For our second result, we first define the Good Algorithmic Regulator (GAR).
Definition 3.1 (Good Algorithmic Regulator, contrastive).
Given the on/off complexities and gap
$a:=K(O^{(N)}_{W,R}),\qquad b:=K(O^{(N)}_{W,\varnothing}),\qquad\Delta:=b-a.$
we say that $R$ is a good algorithmic regulator of gap $\Delta$ for $W$ at horizon $N$ if $\Delta>0$.
Lemma 3.2 (OFF run lower-bounds the world).
There exists $c_{0}=O(1)$ such that
$K\!\big(O^{(N)}_{W,\varnothing}\big)\ \leq\ K(W)+c_{0}\qquad\Rightarrow\qquad K(W)\ \geq\ b-c_{0}.$
Proof.
Given $(W,\varnothing,N)$, the wrapper simulates the OFF dynamics and prints $O^{(N)}_{W,\varnothing}$ with $O(1)$ overhead. ∎
With this definition we can now state and prove our main theorem.
Theorem 3.2 (Probabilistic regulator theorem).
Let $O^{(N)}_{W,R}$ and $\mathsf{E}_{b}^{R}$ be observed and let $\Delta:=K(O^{(N)}_{W,\varnothing})-K(O^{(N)}_{W,R})$. Then there exists $C>0$ such that
$P\!\big((W,R)\mid O^{(N)}_{W,R},\mathsf{E}_{b}^{R}\big)\ \leq\ C\;\cdot\;2^{\,M(W{:}R)}\;2^{-\Delta}.$
Equivalently, every bit by which M(W:R)M(W{:}R) falls short of $\Delta$ costs a factor $\approx 2^{-1}$ in posterior support.
Proof.
(i) Posterior via wrapper. From Eq. (6), $\log_{2}P((W,R)\mid x)\ \leq\ K(x)-K(W,R)+O(1)=a-K(W,R)+O(1).$
(ii) Decompose $K(W,R)$. We use the exact mutual information M(W:R):=K(W)+K(R)K(W,R)M(W{:}R):=K(W)+K(R)-K(W,R), K(W,R)=K(W)+K(R)M(W:R),K(W,R)=K(W)+K(R)-M(W{:}R), hence
K(x)K(W,R)=aK(W)K(R)+M(W:R).K(x)-K(W,R)=a-K(W)-K(R)+M(W{:}R).
(iii) Insert OFF bound (where $b$ enters). By Lemma 3.2, $K(W)\geq b-c_{0}$, so
$K(x)-K(W,R)\ \leq\ M(W{:}R)-(b-a)-K(R)+c_{0}\ =\ M(W{:}R)-\Delta-K(R)+c_{0}.$
(iv) Exponentiate and absorb constants. Exponentiating and using $2^{-K(R)}\leq 1$ gives $P((W,R)\mid x,\mathsf{E}_{b}^{R})\ \leq\ C\,2^{\,M(W{:}R)}\,2^{-\Delta}$ for a constant $C$ absorbing $2^{c_{0}}$ and the wrapper Coding-Theorem constants. ∎
Clarifications.
(i) Where does $b$ appear? Only via Lemma 3.2, which says the OFF run lower-bounds $K(W)$. We never need to compute $b$ explicitly. (ii) Why can we drop $2^{-K(R)}$? A slightly sharper bound is $P((W,R)\mid x,\mathsf{E}_{b}^{R})\leq C\,2^{M(W{:}R)}2^{-\Delta}2^{-K(R)}$. Since $K(R)\geq 0$, dropping $2^{-K(R)}\leq 1$ keeps the focus on the two interpretable scalars $M$ and $\Delta$ without changing the exponential scaling. (iii) Architecture-agnostic. The proof only uses the computable wrapper $(W,R,N)\mapsto x$. Whether $R$ is open- or closed-loop does not affect the posterior algebra. iv) The posterior on the left of Theorem 3.2 is conditioned on the on‑case observation $x$ only. The off‑case run is used solely to supply a numeric lower bound $b:=K(O^{(N)}_{W,\varnothing})$, which implies $K(W)\geq b-O(1)$ by simulation. Formally, we phrase the result as a bound on $\Pr((W,R)\mid x,\mathsf{E}_{b}^{R})$, where $\mathsf{E}_{b}^{R}$ is the side‑event $K(O^{(N)}_{W,\varnothing})=b$”.
As a consequence of Theorem 3.2, one can bound individual posterior masses by $O(2^{K(x)-K(W,R)})$. This implies an exponential tail: $\Pr{M(W:R)\leq\Delta-k}=O(2^{-k})$. In other words, $M(W:R)$ is concentrated within $O(1)$ of its maximum $\Delta$. I.e., there exists $C^{\prime}>0$ (machine/wrapper dependent only) such that for all integers $k\geq 0$,
$\Pr\!\Big\{\,M(W{:}R)\ \leq\ \Delta-k\ \Bigm|\ x,\ \mathsf{E}_{b}^{R}\Big\}\ \leq\ C^{\prime}\,2^{-k}.$
How to read (and use) Theorem 3.2. 1. What we measure: compute the on/off complexities $a=K(O^{(N)}_{W,R})$ and $b=K(O^{(N)}_{W,\varnothing})$ (in practice: fixed MDL code lengths); their difference $\Delta=b-a$ is the compressibility advantage. 2. What the bound says: for any explanation $(W,R)$ of the observed $x$, the universal posterior weight is penalized as $2^{-\Delta}$ unless the pair shares structure: larger M(W:R)M(W{:}R) compensates the penalty. 3. Practical rule of thumb: sustained large $\Delta$ across tasks makes low M(W:R)M(W{:}R) exponentially unlikely. If off‑case $b$ is already small, $\Delta$ will be small—choose a diagnostic readout so the null is not trivially simple.
3.3 Inferring the Objective Function and Planner (As-If Agent)
We next provide a simple theorem regarding the role of complexity as an objective function.
Theorem 3.3 (On/Off evidence equals unconditioned complexity gap).
Under the universal a priori semimeasure,
$\log_{2}\frac{m(O^{(N)}_{W,R})}{m(O^{(N)}_{W,\varnothing})}\;=\;K(O^{(N)}_{W,\varnothing})\;-\;K(O^{(N)}_{W,R})\ \pm O(1).$ (8)
Equivalently, writing the on/off gap as $\Delta:=K(O^{(N)}_{W,\varnothing})-K(O^{(N)}_{W,R})$, we have $m(O^{(N)}_{W,R})/m(O^{(N)}_{W,\varnothing})\;=\;\Theta\!\big(2^{\,\Delta}\big).$ Hence, on the realized pair $(O^{(N)}_{W,R},O^{(N)}_{W,\varnothing})$, maximizing the likelihood of “ON over OFF” is equivalent (up to a constant factor) to minimizing $K(O^{(N)}_{W,R})$ or, equivalently, maximizing the gap $\Delta$.
Proof.
By the Coding Theorem there exist machine-dependent constants $c_{1},c_{2}>0$ such that $c_{1}2^{-K(z)}\leq m(z)\leq c_{2}2^{-K(z)}$ for any string $z$. Apply this to $x$ and $O^{(N)}_{W,\varnothing}$, take base-2 logs, and subtract:
$\log_{2}m(O^{(N)}_{W,R})\;=\;K(O^{(N)}_{W,R})\pm O(1),\qquad-\log_{2}m(O^{(N)}_{W,\varnothing})\;=\;K(O^{(N)}_{W,\varnothing})\pm O(1),$
so $\log_{2}\tfrac{m(O^{(N)}_{W,R})}{m(O^{(N)}_{W,\varnothing})}=K(y)-K(O^{(N)}_{W,R})\pm O(1).$
This statement compares two different strings (the realized ON and OFF outputs) and aligns with the contrastive quantities used elsewhere. The log universal Bayes factor for “ON vs. OFF” is seen to equal the complexity gap $\Delta\pm O(1).$ Thus, on each episode, a regulator behaves as if it were maximizing the scalar $\Delta,$ equivalently minimizing $K\!\big(O^{(N)}_{W,R}\big)$.
Thus, given a regulator $R$ that persistently reduces the readout’s complexity relative to a null baseline $\varnothing$ (the GAR setting of Def. 3.1), we can justify—on purely observational grounds—that $R$ behaves as if it were minimizing a scalar objective. The objective should be canonical (not post hoc) and usable across episodes/tasks.
4 Discussion
We can summarize now our results:
First regulator result: posterior form, given the observed $x$ (Th. 3.1).
By Solomonoff induction and the Coding Theorem [33, 32, 44, 38, 18], we showed that
$\Pr\!\big((W,R)\mid x\big)\;=\;\frac{2^{-K(W,R)+O(1)}}{m(x)}\sim 2^{\,K(x)-K(W,R)}<\frac{1}{\tilde{c}}\,2^{M(W{:}R)}$ (9)
Thus shorter joint generators are exponentially preferred; every extra bit in $K(W,R)$ halves the posterior weight. Decomposing
[…]
shows that, at fixed marginals $K(W),K(R)$, the posterior is exponentially tilted in the algorithmic mutual information M(W:R)M(W{:}R): each extra bit of M(W:R)M(W{:}R) multiplies posterior odds by $\approx 2$.
Second regulator result: posterior with contrast (Th. 3.2).
Without contrast, the story is pure Occam: (9) anchors the posterior near $K(W,R)\approx K(x)$ with a geometric excess-length tail; for fixed $K(W),K(R)$, this yields a high-probability lower bound on M(W:R)M(W{:}R) roughly K(W)+K(R)K(x)K(W)+K(R)-K(x). With contrast, if turning the regulator on yields $K(O^{(N)}_{W,R})=a$ while the off case has $K(O^{(N)}_{W,\varnothing})=b$ with $b>a$, then any explaining $(W,R)$ obeys
$\Pr((W,R)\mid x)\ \leq\ C\,2^{\,M(W{:}R)}\,2^{-\Delta}\!,$
so low mutual information is exponentially disfavored as the gap $\Delta=b-a$ grows. In both regimes, the operational slogan holds: see a simple string ($K(x)$ small), suspect a simple generator ($K(W,R)$ small), and at fixed marginals this means suspect larger M(W:R)M(W{:}R).
The inutition behind these results is that seeing a simple string suggests its generation by a simple program. Formally, for the coupled hypothesis $P=(W,R)$ (wrapped as a single self‑delimiting program), observing $x=O_{W}^{(N)}$ yields the Solomonoff posterior $\Pr(P\mid x)\;\sim 2^{\,K(x)-K(P)},$ by the Coding Theorem [33, 32, 44, 38, 18]. Every extra bit of joint description $K(P)=K(W,R)$ halves posterior weight. This is the quantitative Occam tilt that operationalizes the slogan above.
The posterior mass of joint programs longer than $K(x)+k$ decays geometrically:
$\Pr\{\,K(W,R)\geq K(x)+k\mid x\,\}\ \leq\ 2C\,2^{-k}.$
Hence the typical joint length is near $K(x)$. If $K(W)$ and $K(R)$ are externally constrained (e.g., by design or prior knowledge), this tail translates directly into a lower posterior bound on M(W:R)M(W{:}R) of the form $M(W{:}R)\gtrsim K(W)+K(R)-K(x)-O(\log(1/\delta))$ with posterior confidence $1-\delta$.
Our results are most informative when the observed readout $O_{W}^{(N)}$ is simple. If $K(O_{W}^{(N)})$ is large, the posterior constraints on joint complexity and on mutual information are inherently weak. From the geometric tail, for any $\delta\in(0,1)$ there exists $k=\lceil\log_{2}(2C/\delta)\rceil$ such that, with posterior probability at least $1-\delta$,
$K(W,R)\ \leq\ K(O_{W}^{(N)})+k.$
At fixed marginals $K(W)$ and $K(R)$ this yields
$M(W{:}R)\ \geq\ K(W)+K(R)-K(O_{W}^{(N)})-k-O(\log)\qquad\text{with probability }\geq 1-\delta.$
Hence, if $K(O_{W}^{(N)})$ is large (comparable to $K(W)+K(R)$), the lower bound on M(W:R)M(W{:}R) may be trivial (near 0 up to logs). Intuitively, a complex output does not force shared structure. It is compatible with a complex joint generator even when $W$ and $R$ share little algorithmic information.
On the other hand, the strength of the conclusion depends on the gap $\Delta=b-a$:
$\Pr\!\big((W,R)\mid O_{W}^{(N)},\mathsf{E}_{b}^{R}\big)\ \leq\ C\,2^{\,M(W{:}R)}\,2^{-\Delta},\qquad\Pr\!\big\{M(W{:}R)\leq\Delta-k\ \big|\ O_{W}^{(N)},\mathsf{E}_{b}^{R}\big\}\ \leq\ C^{\prime}2^{-k}.$
Thus even if $a=K(O^{(N)}_{W,R})$ is not very small, a large off/on gap still enforces a large posterior M(W:R)M(W{:}R). In other words, contrast rescues identifiability of shared structure: the evidence scales exponentially in $\Delta$.
In the same universal calculus, regulation carries a canonical scalar interpretation: runtime behavior is as if minimizing $K(O_{W}^{(N)})$ (i.e., maximizing the on/off gap $\Delta$), and design-time comparison across explanations favors larger $M(W{:}R)-\Delta$ via the GAR posterior tilt. This supplies an MDL/Occam objective grounded in the coding theorem (not an ad hoc utility) and complements the IMP’s structural requirements.
We note that a low $K(O_{W}^{(N)})$ alone does not prove high M(W:R)M(W{:}R); it concentrates posterior mass on short joint generators $P$. High M(W:R)M(W{:}R) follows (i) when $K(W)$ and $K(R)$ are fixed/known, or (ii) when contrast pins $K(W)$ high via the off case. Without such constraints, short $P$ could also arise from individually simple $W$ and $R$.
Third regulator result: as-if Objective-function minimization (Th 3.3).
On the realized $O_{W}^{(N)}$, the conditional Coding Theorem gives $\log_{2}\!\big(m(O_{W}^{(N)})/m(O^{(N)}_{W,\varnothing})\big)=K(O^{(N)}_{W,\varnothing})-K(O_{W}^{(N)})$. Thus, the runtime scalar to minimize is $K(O_{W}^{(N)})$. Together with the above, this implies that the regulator is acting (as-if) like an algorithmic agent (with a model of the world, objective function and planner).
Theorem 8 is a representation statement— not a mechanism: $R$ need not compute $K$, but persistent large $\Delta$ is exactly what maximizes universal evidence for “ON”, and it simultaneously makes low M(W:R)M(W{:}R) exponentially unlikely. For a mechanistic objective beyond the Minimum Description Length (MDL) evidence, three constructive routes are standard and complementary. First, in Linear Time-Invariant (LTI) plants the Internal Model Principle makes a structural claim—perfect robust regulation for a specified signal class requires embedding a dynamical copy of the exosystem in the controller—and optimal stabilizing designs arise from explicit quadratic/convex costs (e.g., the Linear Quadratic Regulator, LQR); in the nonlinear case, output‑regulation theory yields constructive regulators under solvable regulator equations together with immersion/detectability and (local) zero‑dynamics stability [12, 13, 34, 20, 17, 25, 3]. Second, in inverse optimal control and inverse reinforcement learning (IRL), trajectories that satisfy Karush–Kuhn–Tucker (KKT) regularity allow identification of a cost $J$ (up to equivalences) whose minimizers reproduce the behavior; in discrete settings, IRL recovers reward functions consistent with observed policies [23, 1, 41]. Third, in revealed‑preference analysis, if cross‑episode choices satisfy the Generalized Axiom of Revealed Preference (GARP), Afriat and Varian guarantee the existence of a strictly increasing, concave utility that rationalizes the data, while Debreu’s representation and the Savage/Karni–Schmeidler frameworks provide (state‑dependent) expected‑utility forms under their axioms [2, 37, 10, 29, 21].
Planner/policy representation (as‑if agent).
Any deterministic causal regulator $R$ induces a computable policy $\pi_{R}:\mathcal{H}_{t}\!\to\!\mathcal{A}$ mapping the coupled history $h_{t}$ (past interface I/O up to time $t$) to the next actuator symbol. This is simply the operational semantics of $R$ viewed as a function of histories.
The coding‑theorem Bayes‑factor identity (Thm. 3.3) supplies a canonical scalar such that, on the realized episode, the sequence of actions produced by $\pi_{R}$ is as if chosen to maximize $J$ subject to the world dynamics. Together with the algorithmic “internal model” conclusion M(W:R)>0M(W{:}R)>0 (i.e., $K(W\mid R)<K(W)$), this yields the standard agent triad:
$\text{(model) }M(W{:}R)>0,\qquad\text{(objective) }J(x)=K(y)-K(x),\qquad\text{(policy/planner) }\pi_{R}.$
Interpretation. This is a representation statement, not a claim that $R$ explicitly solves an optimization problem or contains a modular planner. The existence of $\pi_{R}$ is tautological for any deterministic $R$; the “as‑if” objective follows from the universal evidence identity above. Across tasks/episodes, if the induced choices satisfy standard consistency axioms (e.g., GARP), classical revealed‑preference theorems guarantee the existence of a (monotone, concave) utility that rationalizes the behavior [2, 37]; and in dynamical settings, inverse optimal control / inverse RL constructs a cost for which the observed policy is (near‑)optimal [23, 1]. Thus, given (i) algorithmic model content M(W:R)>0M(W{:}R)>0 and (ii) the canonical scalar $J$ from the coding‑theorem calculus, interpreting the regulator as carrying a policy/planner is both natural and technically justified.
Why AIT is needed
Our results are single-episode and distribution-free: they make statements about an individual realized readout $x$ and about the pair $(W,R)$ as concrete programs, without positing a stochastic source. Classical (Shannon) information theory quantifies expected code lengths and mutual information with respect to a specified probability law; entropy $H(X)$ and mutual information $I(X;Y)$ are undefined without a distribution, and asymptotic statements (AEP/typical sets) further require ergodicity/mixing assumptions (Shannon 1948; Cover–Thomas). In our setting, there is no given probabilistic model over worlds, regulators, or outputs—indeed, the point is to infer model content from a single realized $x$.
AIT supplies exactly the missing calculus. First, it provides a canonical, machine‑invariant complexity for individual strings, $K(x)$, and a universal a priori semimeasure $m(x)$ (Solomonoff–Levin), connected by the Coding Theorem: $\log m(x)=K(x)\pm O(1)$ [33, 32, 43]. This yields a universal Occam posterior over programs, $\Pr\!\big(p\mid x\big)\asymp\ 2^{\,K(x)-|p|},$ from which (i) the geometric excess‑length tail and (ii) our contrastive tilt bounds follow. No analogue exists in Shannon’s framework without positing an external prior over programs; there is no “canonical” $\Pr(p)$ or $\Pr(x)$ in Shannon theory.
Second, AIT lets us formalize “the regulator contains a model of the world” as algorithmic dependence, i.e. positive mutual algorithmic information M(W:R)>0M(W{:}R)>0 (equivalently $K(W\mid R)<K(W)$), a notion defined for individual objects and invariant up to $O(1)$ ([22]). By contrast, Shannon’s $I(W;R)$ requires a joint distribution over $(W,R)$, which is neither given nor natural here.
Third, our key inequalities explicitly use $m(\cdot)$ and prefix complexity: the posterior tilt $2^{\,K(x)-K(W,R)}$, the OFF‑run lower bound on $K(W)$ by simulation, and the contrastive penalty $2^{-\Delta}$ all rely on the Coding Theorem and Kraft–McMillan properties of prefix programs—again, objects absent from Shannon’s ensemble‑level calculus.
Finally, while one can approximate $K(\cdot)$ with MDL/codelengths in practice, MDL’s justification itself rests on the AIT view that shorter descriptions are better and on the coding‑theorem linkage between description length and (universal) probability (Grünwald 2007). In short: AIT provides the universal prior ($m$), object‑level complexities ($K$), and mutual algorithmic information ($M$) needed to turn the informal slogan “see a simple string, suspect a simple generator” into posterior and contrastive theorems—none of which can be stated in Shannon’s framework without ad hoc model classes and priors.
Relation to the Internal Model Principle (IMP)
In the IMP, the closed loop is $(E,C,P)$: an autonomous exosystem $E$ (no inputs and no explicit time dependence, e.g. $\dot{w}=Sw$), a controller $C$ (the regulator), and a plant $P$. The regulated error is e=rye=r-y, where the reference $r$ and disturbances are generated by $E$ and $y$ is measured from $P$ [12, 13]. In our notation, we group the World as $W=(E,P)$ and take the Regulator as $R\equiv C$ (see Figure 2 and Table 1 for the comparison of the two frameworks in the case of a thermostat).
The assumptions in IMP theorems are: (i) Classical necessity is sharpest for finite-dimensional LTI plants (linear, time-invariant) with exogenous signals generated by a finite-dimensional, neutrally stable LTI $E$; stabilizability/detectability and robustness (one fixed $C$ works for a plant neighborhood) are standard [12, 13]. (ii) The structural conclusion is internal-model necessity: perfect robust regulation for the specified signal class requires that $C$ embed a dynamical copy of $E$ (e.g., integrators for steps, oscillators for sinusoids); in MIMO, a $p$-copy is needed. (iii) Nonlinear generalizations (output regulation) require solvability of the regulator equations, suitable immersion/detectability, and (local) stability of the zero dynamics; guarantees are typically local/semiglobal, and necessity is not universal [20, 17, 25]. (iv) Infinite-dimensional/distributed settings and periodic signals may require infinite-dimensional internal models; technicalities arise with unbounded I/O operators [6].
In the AIT formulation (here), we assume: (i) Architecture-agnostic: no required split into $E$ vs. $P$, and no specified place where $R$ enters the causal path; we only assume a computable wrapper mapping $(W,R,N)\mapsto O_{W}$ for a fixed horizon $N$. (ii) Deterministic, closed coupling of world and regulator (no stochastic noise sources into $W$); statements are distribution-free and about the realized sequence. (iii) “Model” means algorithmic dependence: M(W:R)>0M(W\!:\!R)>0 (equivalently $K(W|R)<K(W)$), not a literal dynamical replica. (iv) The main necessity is probabilistic: a positive on/off complexity gap $\Delta=K(O_{W,\emptyset})-K(O_{W,R})$ exponentially tilts the universal posterior against explanations with small M(W:R)M(W\!:\!R); no linearity, smoothness, or regulator-equation conditions are imposed. See Secs. 2–6 of this work.
World (W)Regulator (R)ExosystemPlantMeasurementForcingRefErrorControl
Figure 2: To connect the IMP and the AIT formulation used here, we view the World $W$ as a box containing $E$ and $P$; the Regulator/Controller $R$ (or $C$) is a separate box. Arrows depict Forcing ($E\!\to\!P$), Ref ($E\!\to$ sum), the Error path (sum $\downarrow$ to the world boundary and $\rightarrow R$), and Control ($R\!\to\!P$).
IMP yields a structural necessity (internal model in $C$ of $E$) under explicit dynamical hypotheses; the AIT formulation yields an information‑theoretic necessity (positive M(W:R)M(W\!:\!R) favored by the data) without assuming linearity, an $E/P$ split, or a particular causal insertion point for $R$. The two are complementary: IMP is the backbone for constructive regulation in structured classes; the AIT view covers unstructured architectures and single episodes with a universal Occam calculus [12, 13, 34, 20, 17, 6].
Our statements are thus complementary and distinct: we work in a distribution‑free, program‑level setting and make no linearity or smoothness assumptions. We do not assert the existence of a dynamical replica inside $R$. Instead, we show that sustained contrastive compressibility ($\Delta>0$) tilts the universal posterior toward pairs $(W,R)$ with larger mutual algorithmic information M(W:R)M(W{:}R), i.e., $R$ carries algorithmic structure about $W$. Thus, “the regulator contains a model” is made precise as M(W:R)>0M(W{:}R)>0 (information‑theoretic dependence), not as an embedded exosystem. The IMP supplies structural necessity for perfect regulation within specified signal classes; our AIT results supply information‑theoretic necessity for observed compressibility advantages, beyond linearity or probabilistic assumptions [34].
Role IMP language AIT language (this work) Thermostat instantiation
Exogenous generator Exosystem $E$ = autonomous generator of reference/disturbances; unaffected by $C$. Fold into the World $W$; no architectural split required (you may still conceptually identify this subpart). Reference: setpoint schedule $r(t)$ (often clock‑driven). Disturbances: outdoor temperature, solar load, occupancy heat gains.
Plant Plant $P$ = room thermal dynamics + heater. Also inside World $W$. First‑order building thermal model, heater actuation, heat losses, sensor dynamics.
Controller / Regulator Controller $C$ (the regulator in IMP). Regulator $R$. Thermostat logic: bang‑bang with hysteresis, PI/TPI, or scheduled control.
Measured output $y$. Part of $O_{W}$ (chosen readout). Indoor temperature (or a weighted error signal).
Error / objective e=rye=r-y; IMP concerns asymptotic $e\!\to\!0$ for a signal class. Score regulation by compressibility of a task readout (e.g., $e_{1:T}$) vs. an OFF baseline; gap $\Delta=K(y_{\text{off}})-K(y_{\text{on}})$. Good thermostat $\Rightarrow$ the error stays near a regular pattern (within deadband) $\Rightarrow$ shorter code length than “heater OFF”.
Table 1: Mapping the IMP triple $(E,C,P)$ and the AIT $(W,R)$ formulation to a thermostat. IMP sources: [12, 13, 34]; AIT view: this work.
Practical estimation of $K$ and the gap $\Delta$
The theorems are stated in terms of prefix Kolmogorov complexity, which is not computable. In practice, one fixes a reference prefix code $C$ and estimates
$\widehat{a}:=L_{C}\!\big(O^{(N)}_{W,R}\big),\quad\widehat{b}:=L_{C}\!\big(O^{(N)}_{W,\varnothing}\big),\quad\widehat{\Delta}=\widehat{b}-\widehat{a}$,
with the same compressor $C$ used across all conditions. Persistent $\widehat{\Delta}>0$ across tasks is cumulative evidence that explanations with low M(W:R)M(W{:}R) are exponentially unlikely; maximizing $\widehat{\Delta}$ is the natural scalar objective the regulator appears to optimize on the observed data.
Some standard choices are Lempel-Ziv compressors (LZ77/LZ78/LZW). LZ‑type compressors are universal in a weak sense for stationary ergodic sources and are widely available. Implementations (gzip, lz4, etc.) are practical proxies for $L_{C}(\cdot)$ (Ziv–Lempel 1977; Ziv–Lempel 1978). Recommendations. (i) Fix compressor, window size, and container overhead once; report the raw codelength in bits (not just ratios). (ii) Use identical pre‑processing and record layout across ON/OFF runs. (iii) For short $N$, prefer LZ78‑style (dictionary‑based) or arithmetic‑coded LZ with small headers to reduce fixed overhead. (iv) When comparing across tasks, control for compressor statefulness (reset between episodes).
If both ON and OFF strings are available and one wants a scale‑free sanity check of contrast, compute
$\mathrm{NCD}(x,y)\ :=\ \frac{C(xy)-\min\{C(x),C(y)\}}{\max\{C(x),C(y)\}},$
where $C(\cdot)$ is the chosen code length and $xy$ is concatenation (Cilibrasi–Vitányi 2005). NCD is heuristic but can reveal whether $x$ is “closer” to trivial baselines than $y$.
The Block Decomposition Method (BDM) estimates $K$ by tiling a string (or array) into small blocks whose complexities are looked up from Coding‑Theorem‑Method (CTM) tables (exhaustive output frequency statistics of small machines), plus a logarithmic penalty for multiplicities:
$\widehat{K}_{\mathrm{BDM}}(x)\ \approx\ \sum_{i}K_{\mathrm{CTM}}(b_{i})+\log m_{i},$
where $b_{i}$ are distinct blocks and $m_{i}$ their multiplicities (see [31, 40]). This is sensitive to small‑scale algorithmic regularities beyond LZ’s parse statistics; it works on 1D/2D data (but depends on the chosen CTM table — size and machine model — and it suffers from boundary/tiling effects and additive constants that can be sizable for short $N$).
To improve discrimination, we can i) use paired ON/OFF measurements on the same horizon $N$; report $\widehat{\Delta}$ and its sampling variability across repeats/seeds; ii) include trivial controls (e.g. all‑zero regulator and randomized regulator) to sanity‑check that $\widehat{\Delta}$ responds in the expected direction; iii) for finite $N$, complement point estimates with nonparametric tests (paired permutations on $\widehat{\Delta}$ across episodes); iv) when outputs are multivariate/real‑valued, discretize with a fixed, reported quantization and alphabet before compression.
5 Conclusion
We developed a contrastive, algorithmic formulation of regulation: a regulator $R$ is good for a world $W$ at horizon $N$ when it yields a compressible readout that is strictly more compressible than under a null baseline $\varnothing$. This places the GRT claim (“good regulators are models”) on an AIT footing.
If switching a regulator on makes a system’s measured output much simpler to describe (i.e., more compressible) than when the regulator is off, then the regulator is very likely to carry non‑trivial information about the world it controls—in the precise Algorithmic Information Theory sense of positive mutual algorithmic information between world and regulator. The strength of this evidence grows exponentially with the compressibility gap: large $\Delta$ makes explanations with little shared structure vanishingly likely. Practically, this turns the old cybernetics slogan “every good regulator is a model of the system” into a quantitative, testable claim that does not assume linearity, stochastic models, or specific architectures. On each run, the theorem also singles out a canonical scalar objective: the regulator behaves as if it were minimizing the description length of the realized readout (equivalently, maximizing $\Delta$).
Probabilistically, if $W$ and $R$ are independently sampled minimal programs (no mutual information), then low readout complexity—and especially the contrastive event “low under $R$, high under $\varnothing$”—is exponentially unlikely in $|W|$ and $|R|$. Thus, sustained compressibility relative to baseline is strong evidence that $R$ shares non‑trivial algorithmic structure with $W$ (M(W:R)>0M(W\!:\!R)>0). This is the AIT face of the Good Regulator idea and complements the Internal Model Principle’s structural necessity results for classical regulation: the IMP identifies structural necessities for perfect/robust regulation in classical settings, whereas our AIT view applies beyond linearity and probability and turns regulation into a statement about description length. This bridge clarifies in what limited (yet precise) sense the cybernetics aphorism “good regulators must model” can be made rigorous [8, 5]: successful regulation implies positive mutual algorithmic information between world and regulator.
The result supplies: (i) a distribution‑free, single‑episode diagnostic for “does the controller contain a model?”, (ii) a complement to the IMP (which requires embedding a copy of the signal generator under structured assumptions), and (iii) a simple experimental recipe—fix a lossless compressor, quantize the readout, compute two code lengths (ON vs. OFF), and use their difference $\Delta$ as evidence of model content in the controller.
Finally, the coding‑theorem view identifies a canonical scalar and implicates a planner: runtime minimization of $K(x)$ (equivalently, maximization of $\Delta$).
All together, these results provide the grounds to justify that if a system is seen to regulate another in the algorithmic sense (reducing the complexity of an output of the regulated system compared to no regulation), we can reasonably infer it is likely that the regulator uses a model of the regulated system and an associated scalar objective function.
Acknowledgments
The author thanks Francesca Castaldo for reviewing the manuscript.
Appendix A Appendix
A.1 Setting and core definitions
Universal machine and prefix complexity.
Fix a universal prefix Turing machine $U$. For any finite binary string $x$,
$K(x):=\min\{|p|:U(p)=x\},\qquad m(x):=\sum_{p:\,U(p)=x}2^{-|p|}.$
By the Coding Theorem there exist machine–dependent $c_{1},c_{2}>0$ with $c_{1}2^{-K(x)}\leq m(x)\leq c_{2}2^{-K(x)}$ ( [32]; [43]; Vitányi 2013; Li–Vitányi (book)).
Conditioning convention.
A finite horizon $N\in\mathbb{N}$ is fixed throughout; unless stated otherwise, all complexities are implicitly conditioned on $N$, e.g. $K(x):=K(x\mid N)$ and $m(x):=m(x\mid N)$.
Machines and transcripts.
A world $W$ and regulator $R$ are deterministic causal prefix programs that interact for $N$ steps via interface tapes. Their closed‑loop interaction produces a binary readout $x=O^{(N)}_{W,R}\in\{0,1\}^{N}$. The off/null regulator, denoted $\varnothing$, is the coupling where the regulator’s interface outputs a fixed quiescent symbol (e.g. 0) at all steps, yielding $y=O^{(N)}_{W,\varnothing}$.
Joint description and wrapper.
A fixed constant‑overhead wrapper decodes shortest descriptions of $(W,R)$ and simulates the coupling to print $O^{(N)}_{W,R}$. Denote by $K(W,R)$ the length of a shortest self‑delimiting code for the pair. We use standard chain rules (e.g. $K(W,R)=K(W)+K(R\mid W)\pm O(1)$).
Mutual algorithmic information.
For finite strings $x,y$,
[…]
Equivalently, $M(x{:}y)=K(x)-K(x\mid y)\pm O(\log)$ (Li–Vitányi).
Good Algorithmic Regulator (contrastive).
Let $a:=K(O^{(N)}_{W,R})$ and $b:=K(O^{(N)}_{W,\varnothing})$. Define the gap
$\Delta\ :=\ b-a.$
We say that $R$ is a good algorithmic regulator for $W$ at horizon $N$ if $\Delta>0$. (In practice, $a$ and $b$ are estimated by fixed MDL codelengths; see §4.)
Deterministic upper bound.
Since the wrapper simulates the coupling, one always has
$K\!\big(O^{(N)}_{W,R}\big)\ \leq\ K(W,R)\ \leq\ K(W)+K(R)-M(W{:}R)+O(1).$
A.2 Three-Tape Turing Machine
Definition A.1 (Three-Tape Turing Machine Algorithm).
A three-tape Turing machine algorithm is represented by a Turing machine $T$ with three tapes, and consists of the following components:
  1. 1.
    A finite set of states $Q$, including a designated start state $q_{0}$ and one or more halting states.
  2. 2.
    A finite alphabet $\Sigma$, including a blank symbol, used for the input, output, and private tapes.
  3. 3.
    Three finite tapes, divided into cells, where each cell can contain a symbol from $\Sigma$. These tapes are designated as the input tape, the output tape, and the non-erase private tape.
  4. 4.
    A transition function $\delta:Q\times\Sigma^{3}\rightarrow Q\times\Sigma^{3}\times\{L,R\}^{3}$, defining how the machine moves between states, writes symbols on the three tapes, and moves the tape heads left (L) or right (R) on each tape.
We further identify the state of the private and output states with a set of variables $V=\{v_{1},v_{2},\ldots,v_{n}\}$, with subsets $V_{\text{private}}$ and $V_{\text{output}}$. The time evolution of variables in $V$ is governed by the operation of the Turing machine, as it processes the input, modifies the private tape $V_{\text{private}}$, and writes to the output tape $V_{\text{output}}$, according to $\delta$. So we can also see an algorithm as a specification of the evolution of a set of variables.
The Turing machine begins in the start state with the input written on the input tape and the other tapes blank. It proceeds according to the transition function, writing into the output and private tapes. The private tape can be written to but not erased. When the machine reaches a halting state, the output is read from the output tape.
A.3 Prefix-free programs vs. stop-symbol delimiters (and why it matters)
Setup.
Let $U$ be a universal prefix machine: the domain of its halting programs is prefix-free, so no valid program is a prefix of another. The associated (prefix/self-delimiting) Kolmogorov complexity is
$K_{U}(x)\;=\;\min\{\,|p|\;:\;U(p)=x\text{ and }p\text{ is in a prefix-free domain}\,\}.$
Working with prefix-free domains aligns program lengths with instantaneous (prefix) codes and invokes Kraft–McMillan inequality, the coding-theoretic backbone that underlies many AIT results, including Levin’s universal distribution and the coding theorem [22, 36, 11]. (See also standard IT references for Kraft–McMillan and prefix codes [39, 30].)
Why prefix-freeness is not a mere technicality.
  1. 1.
    Instantaneous decodability and Kraft sums. If the halting programs form a prefix code, then for the multiset of program lengths $\{\,|p|:U(p)\downarrow\,\}$ we have $\sum_{p}2^{-|p|}\leq 1$ by Kraft–McMillan. This lets us interpret $2^{-|p|}$ as a valid “budget” of probability mass per description and leads to semimeasures like Levin’s universal distribution $m_{U}(x)=\sum_{U(p)=x}2^{-|p|}$ with $\sum_{x}m_{U}(x)\leq 1$. This construction is central to algorithmic probability and to the coding theorem (roughly $K(x)\approx-\log m(x)$) [22, 35, 19, 36].
  2. 2.
    Clean invariance and chaining inequalities. The invariance theorem (machine-independence of $K$ up to $O(1)$) and standard chain rules (e.g. $K(x,y)\leq K(x)+K(y\mid x)+O(1)$) are most naturally proved for prefix machines because self-delimitation removes end-of-program ambiguity in compositions and conditional encodings [22, 11].
“Why not just add a stop symbol?”
Suppose we try to avoid the prefix constraint by allowing programs of the form p#p\#, where #\# is an end marker.
  • If the interpreter ignores any trailing bits after #\#, then any extension p#qp\#q yields the same computation as p#p\#. To keep the domain of halting programs unambiguous, you must reject all extensions $p\#q\neq p\#$. But rejecting all such extensions is exactly the prefix-free condition in disguise: no valid codeword is a prefix of another. Thus, a well-implemented “stop-symbol” machine reduces to a prefix-free machine up to a fixed additive overhead for encoding #\#. Consequently, all asymptotic theorems (invariance, coding theorem, bounds using Kraft) remain unchanged up to $O(1)$ [22, 11, 36].
  • If extensions after #\# are allowed as distinct valid programs, then the set of halting inputs is not prefix-free, Kraft–McMillan can fail, and the sum $\sum_{U(p)=x}2^{-|p|}$ need not be bounded by 1. This breaks the semimeasure property essential to Levin’s universal distribution and derails the clean link between probability and description length [35, 19]. In short: allowing arbitrary padding after a nominal “stop” symbol undermines the probability calculus that AIT relies on.
Implications for our results
All conclusions in this paper that rely on (i) the coding-theoretic view of programs, (ii) semimeasures like $m_{U}$, or (iii) standard chain/invariance bounds continue to hold if one uses a stop-symbol formalism implemented so that descriptions are self-delimiting in the sense above. That formalism is equivalent to the prefix-free setting up to $O(1)$ and thus does not change the substance of our arguments or their asymptotic constants. If, however, the stop-symbol scheme admits padded extensions as distinct valid programs, key lemmas using Kraft (and hence bounds derived via $m_{U}$ or coding-theorem arguments) may fail or require nonstandard fixes.
Takeaway.
The “prefix business” is not a dispensable technicality; it encodes self-delimitation that makes programs behave like instantaneous codewords. You can implement self-delimitation via explicit markers, but only if you simultaneously forbid any valid extension after the marker—i.e. you recover a prefix-free domain. With that in place, none of the conclusions elsewhere in the paper need to change (beyond harmless $O(1)$ shifts). Without it, several probability/complexity identifications break.
A.4 Coding Theorems (unconditional and conditional)
Setup and notation.
Fix a universal prefix Turing machine $U$. All logarithms are base 2. For a finite string $x$, let $K(x)$ be its (prefix) Kolmogorov complexity: $K(x):=\min\{|p|:U(p)=x\}$. The universal a priori semimeasure is
$m(x)\;:=\;\sum_{p:\,U(p)=x}2^{-|p|}\,.$
Since the halting programs of a prefix machine form a prefix code, Kraft–McMillan implies $\sum_{x}m(x)\leq 1$.
For conditional versions, we equip $U$ with a read‑only auxiliary input tape that holds side information $y$. Define
$K(x\mid y)\;:=\;\min\{|p|:U(p,y)=x\}\!,\qquad m(x\mid y)\;:=\;\sum_{p:\,U(p,y)=x}2^{-|p|}.$
All $O(1)$ terms and constants below depend only on the choice of $U$, never on $x$ or $y$.
Theorem A.1 (Coding Theorem (unconditional)).
There exist machine‑dependent constants $c_{1},c_{2}>0$ such that for all finite strings $x$,
$c_{1}\,2^{-K(x)}\;\leq\;m(x)\;\leq\;c_{2}\,2^{-K(x)}.$
Equivalently,
$\log m(x)\;=\;K(x)\ \pm O(1).$
Proof sketch.
Lower bound. Let $p^{\star}$ be a shortest program for $x$, so $|p^{\star}|=K(x)$ and $U(p^{\star})=x$. Then $m(x)\geq 2^{-|p^{\star}|}=2^{-K(x)}$ (the constant $c_{1}$ absorbs harmless machine choices).
Upper bound. Because $m(\cdot)$ is a semimeasure, there exists a prefix code with lengths $\ell(x)\leq\lceil-\log m(x)\rceil$ (Shannon–Fano/Kraft–McMillan). A fixed decoder transforms the codeword for $x$ into $x$, so $K(x)\leq\ell(x)+O(1)\leq-\log m(x)+O(1)$. Rearranging gives $m(x)\leq c_{2}\,2^{-K(x)}$. ∎
Theorem A.2 (Coding Theorem (conditional)).
There exist machine‑dependent constants $c^{\prime}_{1},c^{\prime}_{2}>0$ such that for all finite strings $x,y$,
$c^{\prime}_{1}\,2^{-K(x\mid y)}\;\leq\;m(x\mid y)\;\leq\;c^{\prime}_{2}\,2^{-K(x\mid y)}.$
Equivalently,
$\log m(x\mid y)\;=\;K(x\mid y)\ \pm O(1).$
Proof sketch.
Lower bound. With $p^{\star}$ a shortest conditional program for $x$ given $y$, we have $U(p^{\star},y)=x$, hence $m(x\mid y)\geq 2^{-|p^{\star}|}=2^{-K(x\mid y)}$.
Upper bound. For fixed $y$, $m(\cdot\mid y)$ is a semimeasure, so there is a prefix code (depending on $y$) with $\ell(x\mid y)\leq\lceil-\log m(x\mid y)\rceil$ and a fixed decoder (shared across all $y$) that maps codewords plus $y$ to $x$. Therefore $K(x\mid y)\leq-\log m(x\mid y)+O(1)$, which rearranges to the stated upper bound. ∎
Remarks.
  • The constants $c_{1},c_{2},c^{\prime}_{1},c^{\prime}_{2}$ (and all $O(1)$ slacks) depend only on the choice of the universal prefix machine $U$; changing $U$ shifts $K(\cdot)$ by at most an additive constant (invariance theorem), which becomes a multiplicative constant on $m(\cdot)$.
  • Theorems A.1A.2 are often summarized as $m(x)\asymp 2^{-K(x)}$ and $m(x\mid y)\asymp 2^{-K(x\mid y)}$, read “within constant factors”.
  • Immediate corollaries used in the main text include the posterior under the universal prior: for any program $p$ with $U(p)=x$,
    $\Pr\{p\mid x\}=\frac{2^{-|p|}}{m(x)}\in\Big[\tfrac{1}{c_{2}},\tfrac{1}{c_{1}}\Big]\!\cdot 2^{\,K(x)-|p|},$
    and the geometric excess‑length tail: $\Pr\{|p|\geq K(x)+k\mid x\}\leq C\,2^{-k}$ for some constant $C>0$.
A.6 Single-episode compressibility is non-diagnostic
Intuitively, knowing that the regulator-world coupled system produces a low-complexity world output $x$ reduces the set of possible worlds to select from. In turn, this allows for a shorter description of the world using $R$ and the complexity bound of the output. The program may say: “To specify $W$, run the dynamics for all possible $W$-$R$ pairs and delete all world model candidates with complex outputs (above the set complexity bound $K(x)<a$). Then use a reduced index to identify WW”. This means that K(W|R,K(x)<a)<K(W)K(W|R,``K(x)<a")<K(W), which implies M(W;R|K(x)<a)>0M(W;R|``K(x)<a")>0.
Theorem A.4.
(low complexity output $\Rightarrow$ strict but tiny shrinkage) Fix a universal prefix machine. Let $m:=|W|$ and $r:=|R|$ denote minimal code lengths, and let $N\geq m$. For a fixed regulator $R$ and horizon $N$, consider the class $\mathcal{P}_{m}$ of all minimal $m$-bit world programs. Assume we only know that the closed-loop transcript has low complexity,
$E_{a}:\qquad K\!\big(O^{(N)}_{W,R}\mid R,N\big)\ \leq\ a,$
for some threshold a<mca<m-c, where $c=O(1)$ is a machine-dependent constant. Claim. The set of candidates consistent with $E_{a}$ is a strict subset of $\mathcal{P}_{m}$:
$S_{R,N,a}(m)\ :=\ \Big\{\,W\in\mathcal{P}_{m}\ :\ K\!\big(O^{(N)}_{W,R}\mid R,N\big)\leq a\,\Big\}\ \subsetneq\ \mathcal{P}_{m}.$
Consequently,
$K\!\big(W\mid R,E_{a}\big)\ \leq\ \log_{2}\!\big(|\mathcal{P}_{m}|-1\big)\ <\ \log_{2}|\mathcal{P}_{m}|\ =\ m\ \pm O(1),$
i.e., strictly $K(W\mid R,E_{a})<K(W)$ (by a vanishingly small amount).
Proof. By Kleene’s recursion theorem (quines), there exists a program $W^{\star}\in\mathcal{P}_{m}$ that prints its own source as the first $m$ output bits and then halts (or pads). Hence $K\!\big(O^{(N)}_{W^{\star},R}\mid R,N\big)\geq K(W^{\star})-O(1)=m-O(1)>a$, so $W^{\star}\notin S_{R,N,a}(m)$. Therefore $S_{R,N,a}(m)\subsetneq\mathcal{P}_{m}$, implying $\log|S_{R,N,a}(m)|<\log|\mathcal{P}_{m}|=m\pm O(1)$. $\square$
To see how small the information gained can be, consider a world program $W$ whose last line is “print $u\times O_{R}$,” where $u$ is some computed world variable. If $R$ simply outputs 0, the world output becomes the all‑zeros string, hence very compressible. Knowing that $R$ outputs 0 and that the world output is $0^{N}$ does restrict the structure of the world program (it must include the final multiplication by the regulator output, or something similar on the realized trace), but that restriction can be tiny—the calculation of $u$ may still be arbitrarily complex.
Although we have shown that $R$ and $E_{a}$ together share information with $W$, it may be very small for any given case, and, in any case, this does not imply that $R$ and $W$ share information. The chain rule gives
$M\big(W:(R,E)\big)$ $=K(W)+K(R,E)-K(W,R,E)$
$=K(W)+\big[K(R)+K(E\mid R)\big]-\big[K(R)+K(W,E\mid R)\big]\pm O(\log)$
$=\underbrace{K(W)+K(R)-K(W,R)}_{M(W:R)}\;+\;\underbrace{K(W\mid R)+K(E\mid R)-K(W,E\mid R)}_{M(W:E\mid R)}\;\pm O(\log).$
Thus, knowing that the coupled $(W,R)$ system produces a low-complexity readout $x$ in a single run strictly prunes the set of candidate worlds, but in the worst case this shrinkage is only $O(1)$ and—critically—does not by itself imply M(W:R)>0M(W{:}R)>0; it certifies at most $M\!\big(W:(R,E_{a})\big)>0$ via the chain rule.
Does contrast fix the non‑probabilistic identifiability? Let $E_{a,b}$ be the (contrastive) event
$E_{a,b}:\qquad K\!\big(O^{(N)}_{W,R}\big)\leq a\quad\text{and}\quad K\!\big(O^{(N)}_{W,\varnothing}\big)\geq b\ (b>a).$
The deterministic shrinkage equals
$K(W)-K\big(W\mid R,E_{a,b}\big)\;=\;M\!\big(W:(R,E_{a,b})\big)\ \pm O(\log),$
and by the chain rule this splits as
$M\!\big(W:(R,E_{a,b})\big)\;=\;M(W{:}R)\;+\;M\!\big(W{:}E_{a,b}\mid R\big)\ \pm O(\log).$ (12)
Thus, from single‑episode ON/OFF facts we can certify at most $M\!\big(W:(R,E_{a,b})\big)>0$; in general this does not imply M(W:R)>0M(W{:}R)>0, because the conditional term $M(W{:}E_{a,b}\mid R)$ can carry (almost) all the gain or because of synergy.
Furthermore, even if the mutual algorithmic information between world and regulator is null, it may be the case that coupling them leads to a reduction of complexity in the world output by chance.
These caveats motivate the probabilistic analysis in the paper.
We discuss in more detail the case of synergy, and also show that a decrease of complexity cannot certify mutual information in a particular case.
Chain rule and a synergy counterexample.
By the chain rule for mutual information,
$M\!\big(W:(R,E_{a,b})\big)\;=\;M(W{:}R)\;+\;M\big(W{:}E_{a,b}\mid R^{\*}\big)\;+\;O(\log n),$ (13)
where $R^{\*}$ is a shortest description of $R$ (drop $R^{\*}$ and the $O(\log n)$ term in the Shannon case). [^2] Algorithmic version: Li & Vitányi, An Introduction to Kolmogorov Complexity and Its Applications, 4th ed., Springer, 2019. Shannon version: Cover & Thomas, Elements of Information Theory, 2nd ed., Wiley, 2006. Thus, observing that $M\!\big(W:(R,E_{a,b})\big)>0$ does not imply M(W:R)>0M(W{:}R)>0, because the conditional term $M\big(W{:}E_{a,b}\mid R\big)$ can carry (almost) all of the gain.
Example (XOR/synergy). Let $R,E_{a,b}\in\{0,1\}^{n}$ be independent, incompressible strings, and set $W=R\oplus E_{a,b}$ (bitwise XOR). Then:
$M(W{:}R)$ $\stackrel{{\scriptstyle+}}{{=}}K(W)+K(R)-K(W,R)$
$\stackrel{{\scriptstyle+}}{{\leq}}K(W)-K(W\mid R)$
$\stackrel{{\scriptstyle+}}{{=}}K(W)-K(E_{a,b}\mid R)$
$\stackrel{{\scriptstyle+}}{{\leq}}O(\log n),$ (14)
because $E_{a,b}\mapsto W$ is a bijection given $R$ and independence gives $K(E_{a,b}\mid R)\stackrel{{\scriptstyle+}}{{\geq}}n$. In contrast,
$M\!\big(W:(R,E_{a,b})\big)$ $\stackrel{{\scriptstyle+}}{{=}}K(W)-K\!\big(W\mid R,E_{a,b}^{\*}\big)$
$\stackrel{{\scriptstyle+}}{{\geq}}n-O(\log n),$ (15)
since $K(W\mid R,E_{a,b})=O(1)$ and $K(W)\stackrel{{\scriptstyle+}}{{\geq}}K(W\mid R)\stackrel{{\scriptstyle+}}{{\geq}}n$. Hence the conditional term $M\big(W{:}E_{a,b}\mid R\big)$ carries essentially all the information. For the Shannon analogue, take $R,E_{a,b}\sim\mathrm{Ber}(\tfrac{1}{2})$ i.i.d.; then $I(W;R)=0$, $I(W;E_{a,b}\mid R)=H(W)=n$, so $I\!\big(W;(R,E_{a,b})\big)=n$. [^3] XOR–synergy as a canonical case in multivariate information: Williams & Beer (2010). For the identity $M(x{:}y)=K(x)-K(x\mid y^{\*})+O(\log)$ used above, see Bennett, Gács, Li, Vitányi & Zurek, IEEE Trans. Inf. Theory, 1998.
Chance simplification with $M(W{:}R)\approx 0$ is possible.
Fix a universal prefix Turing machine $U$ and a finite horizon $N$. All complexities are implicitly conditioned on $N$ (we write $K(\cdot)$ for $K(\cdot\,|\,N)$). Identify Turing machines with their shortest prefix codes and write $|W|=K(W)$, $|R|=K(R)$. The coupled world–regulator system produces a deterministic readout
$x\;:=\;O^{(N)}_{W,R}\in\{0,1\}^{N}.$
There is no auxiliary map: a fixed, constant‑overhead wrapper decodes $(W,R)$ and simulates the interaction to print $x$ (decode+simulate). Consequently
$K(x)\;\leq\;K(W,R)+O(1)\;=\;K(W)+K(R)-M(W{:}R)\pm O(\log N),$ (16)
and we use the standard identity $M(X{:}Y)=K(X)-K(X\mid Y^{\*})\pm O(\log)$. (See eq. (1) and the chain‑rule algebra in §2–3 of the WP.) [^4] For textbook background on prefix complexity, chain rules, and $M(x{:}y)=K(x)-K(x\mid y^{\*})\pm O(\log)$, see Li & Vitányi (2019), and Bennett et al. (1998).
For concreteness in the examples below we take $|W|=|R|=n$ and set $N=n$; this is only for clarity (all statements have the obvious adjustments if $N\neq n$).
Claim (It can happen that $K(x)$ is small while $M(W{:}R)\approx 0$).
There exist pairs $(W,R)$ with $M(W{:}R)=O(\log n)$ such that the coupled output $x=O^{(N)}_{W,R}$ has very small complexity (e.g. $K(x)=O(\log N)$).
Construction (existence, uses only the W+RW{+}R coupling). Fix a threshold $\Delta\in\{1,\dots,N\}$. Define a world program $W_{\Delta}$ that monitors the first $\Delta$ symbols emitted by the regulator on the interface and then latches:
[…]
Here $z$ is hard‑coded in $W_{\Delta}$ (so $K(z)\stackrel{{\scriptstyle+}}{{=}}N$ and $K(W_{\Delta})\stackrel{{\scriptstyle+}}{{=}}|W|=n$). Choose any regulator $R^{(\Delta)}$ whose first $\Delta$ interface outputs are $0^{\Delta}$ and whose remaining behavior is generated by a shortest program of length $\stackrel{{\scriptstyle+}}{{=}}n$ independent of $W_{\Delta}$. Then
$M(W_{\Delta}{:}R^{(\Delta)})=O(\log n)\qquad\text{but}\qquad x=0^{N}\;\Rightarrow\;K(x)=O(\log N).$
Thus, even with $M(W{:}R)\approx 0$ (up to the usual $O(\log)$ slack), the coupled program can, on the realized episode, yield a low‑complexity output.
“Rare but possible” bound (balanced couplings).
Suppose the world implements a balanced dependence on the regulator’s interface in the sense that, for fixed $W$, the map $u\mapsto x$ is a permutation of $\{0,1\}^{N}$ when we view $u:=O_{R}[1{:}N]$ as the regulator’s output sequence (e.g., the world computes $x=z\oplus u$ with a fixed $z=z(W)$). If $R$ is sampled independently and its interface sequence $u$ is (close to) uniform on $\{0,1\}^{N}$ (e.g., drawn from a family with pseudorandom outputs), then by the standard Kolmogorov counting bound (at most $2^{k+1}$ NN‑bit strings have $K\leq k$),
$\Pr\big[K(x)\leq k\big]\;\leq\;2^{\,k+1-N}.$
Equivalently, the probability of a $\Delta$‑bit drop ($K(x)\leq N-\Delta$) is $\leq 2^{\,1-\Delta}$. Thus, a very simple $x$ can occur by chance, but only with exponentially small probability in the amount of simplification. [^5] Counting bound: at most $2^{k+1}$ strings of length $N$ have complexity $\leq k$; see Li & Vitányi (2019).
Ex‑post constraint when $R$ is invertible from $(W,x)$.
If the coupled architecture allows recovery of $R$ from $(W,x)$ via a computable inverse (i.e., there exists a fixed decoder such that $R=G(W,x)$), then
$K(x)\;\geq\;K(R\mid W^{\*})-O(1)\;=\;K(R)-M(W{:}R)-O(\log n).$
Hence, with $K(R)\stackrel{{\scriptstyle+}}{{=}}n$ and $M(W{:}R)\approx 0$, a large drop in $K(x)$ cannot occur under such invertible (in $R$) couplings. When a very small $x$ is observed in this case, it forces M(W:R)M(W{:}R) to be large. (Identity used: $M(X{:}Y)=K(X)-K(X\mid Y^{\*})\pm O(\log)$.)